a paper clip is twelve over sixteen inches long.If you measured in One over four inches,what would the measurement be?

Answer by stanbon(75887)

a paper clip is twelve over sixteen inches long.If you measured in One over four inches,what would the measurement be?
----------------------
(12/16)/(1/4) = (12/16)*(4/1) = (3/4)*(4/1) = 3
================
Cheers,
Stan H.